# Generation of complete graph states in a spin-$1/2$ Heisenberg chain with a globally optimized magnetic field

### 07 Jan 2024 v2

@@ Abstract
Graph states possess significant practical value in measurement-based quantum computation, with complete graph states that exhibit exceptional performance in quantum metrology. In this work, we introduce a method for generating multiparticle complete graph states using a spin-$1/2$ Heisenberg $XX$ chain subjected to a time-varying magnetic field, which applies to a wide range of systems. Our scheme relies exclusively on nearest-neighbor interactions between atoms, with real-time magnetic field formation facilitated by quantum optimal control theory. We focus specifically on neutral-atom systems, finding that multiparticle complete graph states with $N=3\sim 6$ can be achieved in less than $0.25~{}\mu{\rm s}$, utilizing a hopping amplitude of ${J}/{(2\pi)}=-2.443~{}{\rm MHz}$. This assumes an initial state provided by an equal-weight superposition of all spin states that are encoded by the dipolar interacting Rydberg states. Additionally, we thoroughly address various experimental imperfections and showcase the robustness of our approach against atomic vibrations, fluctuations in pulse amplitude, and spontaneous emission of Rydberg states. Considering the common occurrence of disturbances in experimental setups of neutral-atom systems, our one-step strategy for achieving such graph states emerges as a more empirically viable alternative to techniques based on controlled-Z gates.
@@ I Introduction
Graph states, derived from mathematical graph theory Hein et al. (2004); Wu et al. (2014), play a pivotal role in quantum information theory. It describes a class of multiparticle-entangled states including Greenberger-Horne-Zeilinger (GHZ) states and cluster states Briegel and Raussendorf (2001). Such states form a universal resource for measurement-based quantum computing (MBQC) without any controlled two-system quantum gates Raussendorf and Briegel (2001); Bacon and Flammia (2010); Gao et al. (2011); Menicucci (2014); Reimer et al. (2019). Some of which can also be widely applied in quantum teleportation Ren et al. (2008); Zhou et al. (2019); Kazemikhah et al. (2022) and quantum secret sharing Markham and Sanders (2008); Keet et al. (2010). Multiparticle complete graph states represent that any two particles interact with each other, requiring more projective measurements to be disentangled. By performing a single $\sigma_{y}$ measurement at an arbitrary vertex (qubit), these states can be converted to star graph states (GHZ type) Hein et al. (2004); Van den Nest et al. (2004). Such states exhibit excellent performance in quantum metrology Tóth and Apellaniz (2014); Shettell and Markham (2020), quantum error-correction encoding Liao et al. (2022), and quantum cryptography Unnikrishnan and Markham (2022). However, although complete graph states have many advantages, less attention has been paid to preparing such graph states experimentally, focusing mainly on linear and 2D cluster states Schwartz et al. (2016); Reimer et al. (2019); Asavanant et al. (2019); Yang et al. (2022a).
@@
The preparation of multiparticle complete graph states can be conceptually divided into two primary categories. The first approach is applicable to arbitrary graph states, employing the mathematically proven two-qubit controlled-Z gate (CZ) Cho and Lee (2005); Lin et al. (2008); Ananth and Senthilvelan (2016). However, these protocols are constrained by the operational time and fidelity of the logic gate. Conversely, multiparticle complete graph states can also be generated within the framework of cavity quantum electrodynamics Schön et al. (2005); Zhang et al. (2007); Gonţa et al. (2009). In particular, certain theoretical schemes Li et al. (2008, 2008); Shao and Zhang (2008) enable a one-step realization, but these one-step approaches require uniform interaction strengths between all particle pairs, which poses challenges for experimental implementation. Recently, the proposal of an equivalent Hamiltonian (Heisenberg $XXX$ model with a staggered field) Gietka et al. (2021) has sparked interest in Floquet engineering for constructing such a Hamiltonian Geier et al. (2021); Scholl et al. (2022). Nevertheless, in practice, controlling the impact of noncommutation of the Hamiltonian under periodic driving becomes challenging as the particle number increases.
@@
As quantum control theory progresses, optimal quantum control furnishes a toolkit for crafting external field shapes to fulfill assigned tasks, such as state generation Fisher et al. (2009); Glaser et al. (2015); Lovecchio et al. (2016); Omran et al. (2019), quantum gate implementation Goerz et al. (2014); Propson et al. (2022), and noisy intermediate-scale quantum technology Preskill (2018). These tools aim to achieve the specified objectives with minimal energy and resource expenditure. The practical realization of calculated field shapes has been facilitated by the availability of commercially accessible arbitrary waveform generators. In this paper, leveraging the Gradient Ascent Pulse Engineering (GRAPE) optimization algorithm Khaneja et al. (2005); Wu et al. (2019), we propose an alternative Hamiltonian represented as a spin-$1/2$ Heisenberg $XX$ chain with identical nearest-neighbor interactions under an optimized time-dependent magnetic field. Governed by this Hamiltonian, we successfully generate complete graph states for a particle number ranging from $N=3$ to $N=6$. Our protocol boasts two key advantages: (i) It exclusively relies on the nearest-neighbor interactions, ensuring robust operability and scalability in experimental setups. (ii) The operational time is significantly reduced to less than $0.25~{}\mu{\rm s}$, a crucial factor in mitigating decoherence effects.
@@
The structure of this paper is organized as follows. In Sec. II, we initiate our exploration with the introduction of the complete graph state’s definition and an analysis of the corresponding scheme. Subsequently, in Sec. III, we employ the GRAPE optimization algorithm to articulate a system-independent realization of complete graph states in a “one-step” fashion. Transitioning to Sec. IV, we delve into a specific physical system—the Rydberg atom array—providing a comprehensive examination of the experiment’s feasibility. Our results suggest that accounting for experimental noise, the population of the complete graph state with $N=6$ can exceed $0.93$ within a time frame of $0.25~{}\mu{\rm s}$. Finally, Sec. V encapsulates our findings, offering a succinct summary.
@@
Figure 1: (a) Multiparticle complete graph states with the number of vertices $N=3\sim 6$. The graph connecting the two associated vertices has an edge whenever two particles have interacted. (b) Experimental geometry. The particles are arranged as a chain with the open boundary condition and have equivalent interactions between the nearest particles.
@@
with the convention $\sigma_{z}^{N+1}\equiv 1$. To generate such complete graph states, we construct a Hamiltonian holding a global time-depending magnetic field, which takes the form of ($\hbar=1$)
@@
$H_{\textrm{XX}}=\sum_{\langle i,j\rangle}J(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j})$ (3)
@@
Since the two parts ($H_{\textrm{XX}}$ and the global magnetic field term) commute with each other, the evolution can take place under the Hamiltonian (2) or under the $H_{\textrm{XX}}$ and the global magnetic field separately. Theoretically, any equivalent $H_{\textrm{XX}}$ model can be applied to this scenario. Under the global magnetic field with constant amplitude, an exact analytical solution of $B$ can be found for $N=3$, as shown in Appendix A. However, for $N>3$, such an exact solution is nonexistent. Thus, we consider a time-dependent magnetic field $B(t)$ and introduce the GRAPE optimization algorithm to find the optimal solution for the system. In this way, we can expect to obtain a multiparticle universal protocol.
@@ III The utilization of the GRAPE optimization algorithm
Given the known form of the target state $|K_{N}\rangle$, the optimization problem, which is implemented using the gradient descent algorithm, is translated into the search for optimal amplitudes $B(t)$ of the magnetic field. These amplitudes aim to guide the initial state $\rho(0)$ to $\rho(T)$ with maximum similarity to $\rho_{K}=|K_{N}\rangle\langle K_{N}|$ within a specified time duration $T$. In our protocol, the system Hamiltonian is segmented into two distinct parts.
@@
where $H_{z}=\sum_{i=1}^{N}S^{z}_{i}$ is referred to as the control component. The quantum control landscape, depicting the expectation value of a Hermitian observable operator $\rho_{K}$ at a given time $T$, is represented as HO2006226
@@
Target state Evolution time ($J^{-1}$) Population
@@
The results demonstrate the successful generation of multiparticle complete graph states under the influence of a spin-$1/2$ Heisenberg $XX$ chain in a time-varying magnetic field optimized by the GRAPE algorithm. Consequently, the preparation of multiparticle complete graph states can be achieved mathematically in a “one-step” manner through Eq. (2). This protocol is universally applicable in systems capable of nearest-neighbor interactions characteristic of the spin-$1/2$ Heisenberg $XX$ chain, including superconducting quantum circuits Zippilli et al. (2015); Roushan et al. (2017a, b); Yanay et al. (2020) and ion traps Graß and Lewenstein (2014); Brydges et al. (2019); Davoudi et al. (2020); Birnkammer et al. (2022); Kranzl et al. (2023). In the following part, we will delve into the detailed methodology for preparing the aforementioned multiparticle complete graph states within a neutral-atom system, closely considering experimental parameters.
@@ IV A specific implementation based on Rydberg atom array
Figure 3: The realization of the multiparticle complete graph states with $N=3\sim 6$ governed by the Hamiltonian $H_{\textrm{sys}}$ initialized from state $|\Psi_{0}\rangle=\bigotimes_{i=1}^{N}[{1}/{\sqrt{2}}(|\uparrow\rangle+|\downarrow\rangle)]$. (a) and (b) respectively show evolution under the magnetic field in Gaussian and random types for $N=3$. The dotted lines correspond to the evolution under the guess field, while the solid lines correspond to the optimized one. Then (c) and (d) show the evolution for $N=4$, while (e) and (f) for $N=5$, (g) and (h) for $N=6$. The insets show the corresponding time-dependent $B(t)$ of the initial guess and the optimized result, where $B_{0}=2\pi\times 1~{}{\rm MHz}$.
The Rydberg atom array, characterized by robust programmability, has recently emerged as a powerful candidate in quantum simulations, spurred by breakthroughs in experimental technology. Individual Rydberg atoms, for example, can be put into an optical tweezer array Endres et al. (2016); Barredo et al. (2018). Leveraging the long-range interactions inherent in highly excited Rydberg states, the Rydberg atom array finds widespread applications in quantum state transmission Buchmann et al. (2017), quantum topology research de Léséleuc et al. (2019); Scholl et al. (2021); Giudici et al. (2022), and artificial gauge field Celi et al. (2014); Yang et al. (2022b). Notably, recent works, such as those in Refs. Ebadi et al. (2021); Bluvstein et al. (2022), have realized a primary quantum computer employing 256 Rydberg atoms, demonstrating proficiency in various quantum tasks.
Within the Rydberg atom array, the Rydberg dipole-dipole interaction serves as a direct means to achieve $H_{\textrm{XX}}$. Additionally, the magnetic field $B(t)$ can be equivalently obtained through the Stark effect induced by a large detuned time-dependent laser field coupled transition between the intermediate state $|p\rangle$ and one of the Rydberg states. It is crucial to note that the inevitable long-range interactions among Rydberg atoms, such as van der Waals (vdW) forces and the next-nearest-neighbors (NNN) dipolar interactions, should also be taken into consideration, regarding to the realistic physical system.
@@ IV.1 The realization of multiparticle complete graph states
We consider a one-dimensional chain geometry with open periodic boundary conditions, as depicted in Fig. 1(b), in which $N$ atoms are contained in tweezers that are equidistantly spaced $R$ along the quantized $z$ axis. For the purpose of illustration, states with opposite parity are transformed into pseudospin states, using ${}^{87}\textrm{Rb}$ atom as an example, $|\uparrow\rangle=|80S_{1/2},m_{j}=1/2\rangle$ and $|\downarrow\rangle=|79P_{3/2},m_{j}=1/2\rangle$ Browaeys et al. (2016). The resonant dipole-dipole interaction gives rise to the Heisenberg $XX$ Hamiltonian
$H_{\textrm{XX}}=\sum_{\langle i,j\rangle}V_{\textrm{dip}}(\theta)(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}),$ (12)
where $V_{\textrm{dip}}(\theta)=C_{3}(1-3\cos^{2}\theta)/R^{3}$ with $\theta$ the polar angle between the quantization $z$ axis and the vector direction of atomic connection, and the constant $C_{3}$ is defined as $2\pi\times 8.780~{}{\rm GHz}\cdot\mu{\rm m}^{3}$ S̆ibalić et al. (2017). For $\theta=0$ and $R=19.3~{}\mu{\rm m}$, we have $V_{\textrm{dip}}\simeq-2\pi\times 2.443~{}{\rm MHz}$. Additionally, we account for inevitable interactions, including the vdW interaction and the NNN dipolar interactions, which are characterized as error terms
$H_{\textrm{err}}$ $=$ $\sum_{\langle i,j\rangle}(U^{\uparrow}_{ij}|\uparrow_{i}\uparrow_{j}\rangle\langle\uparrow_{i}\uparrow_{j}|+U^{\downarrow}_{ij}|\downarrow_{i}\downarrow_{j}\rangle\langle\downarrow_{i}\downarrow_{j}|)$ (13)
$+\sum_{\langle\langle i,j\rangle\rangle}\frac{V_{\textrm{dip}}}{8}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}),$
where $U^{\uparrow(\downarrow)}_{ij}=-C^{\uparrow(\downarrow)}_{6}/R^{6}$ with $C^{\uparrow}_{6}=-2\pi\times 4161.55~{}{\rm GHz}\cdot\mu{\rm m}^{6}$ and $C^{\downarrow}_{6}=2\pi\times 3452.60~{}{\rm GHz}\cdot\mu{\rm m}^{6}$. $\langle\langle i,j\rangle\rangle$ indicate the summation over the next-nearest neighbors. Thus, the system Hamiltonian can be rewritten as
@@
Using $N=3\sim 6$ as illustrative examples, we depict the evolutions governed by Eq. (14). In Fig. 3, we specifically show the evolutions under the initial guess and optimized fields in Gaussian and random types. Within this framework, the populations in the target state, denoted as $P(N)_{O}$, exhibit multiple peak points corresponding to different values of $T$ (as shown in the picture in Appendix B). By scanning the time within $0.25~{}\mu{\rm s}$, we successfully obtain the optimized fields under $H_{\textrm{sys}}$ for various numbers of particles. The inset shows the time-dependent $B(t)$ of the application, where solid lines represent the optimized pulse, and dotted lines represent the guess pulse. The evolution time $T$ and the populations of the target states are summarized in Table 2. Although the population of the target state decreases as $N$ increasing, we can still prepare the six-particle complete state with population over $0.94$.
@@
$|K_{3}\rangle$ 0.141 0.9988
$|K_{4}\rangle$ 0.172 0.9911
$|K_{5}\rangle$ 0.205 0.9781
$|K_{6}\rangle$ 0.237 0.9443
Recently, a quantum circuit utilizing the Rydberg atom array has been demonstrated to generate a complete graph state of five particles Crescimanna et al. (2023). When coupled with quantum annealing algorithms, the fidelity of the target state can be enhanced from approximately $0.85$ to $0.999$ by increasing the computational depth. However, this heightened computational depth is accompanied by prolonged execution times, leading to an increased susceptibility to decoherence effects. In contrast, our protocol enables the generation of the target state in a single step, regardless of the number of atoms involved. This approach significantly reduces operational time while preserving high levels of fidelity, presenting notable advantages in mitigating decoherence effects.

### 11 Mar 2024 v3

@@ Abstract
Graph states possess significant practical value in measurement-based quantum computation, with complete graph states that exhibit exceptional performance in quantum metrology. In this work, we introduce a method for generating multiparticle complete graph states using a spin-$1/2$ Heisenberg $XX$ chain subjected to a time-varying magnetic field, which applies to a wide range of systems. Our scheme relies exclusively on nearest-neighbor interactions between atoms, with real-time magnetic field formation facilitated by quantum optimal control theory. We focus specifically on neutral-atom systems, finding that multiparticle complete graph states with $N=3\sim 6$ can be achieved in less than $0.25~{}\mu{\rm s}$, using a hopping amplitude of ${J}/{(2\pi)}=-2.443~{}{\rm MHz}$. This assumes an initial state provided by an equal-weight superposition of all spin states that are encoded by the dipolar-interacting Rydberg states. Furthermore, we thoroughly address various experimental imperfections and showcase the robustness of our approach against atomic vibrations, fluctuations in pulse amplitude, and spontaneous emission of Rydberg states. Taking into account the common occurrence of disturbances in the experimental setups of neutral-atom systems, our one-step strategy to achieve such graph states emerges as a viable alternative to techniques based on controlled-Z gates.
@@ I Introduction
Graph states, derived from mathematical graph theory [1, 2], play a central role in quantum information theory. It describes a class of multiparticle-entangled states including Greenberger-Horne-Zeilinger (GHZ) states and cluster states [3]. Such states form a universal resource for measurement-based quantum computing (MBQC) without controlled two-system quantum gates [4, 5, 6, 7, 8]. Some of these can also be widely applied in quantum teleportation [9, 10, 11] and quantum secret sharing [12, 13]. Multiparticle complete graph states represent that any two particles interact with each other, requiring more projective measurements to be disentangled. When a single $\sigma_{y}$ measurement is performed at an arbitrary vertex (qubit), these states can be converted to star graph states (GHZ type) [1, 14]. Such states exhibit excellent performance in quantum metrology [15, 16], quantum error-correction encoding [17], and quantum cryptography [18]. However, although complete graph states have many advantages, less attention has been paid to preparing such graph states experimentally, focusing mainly on linear and 2D cluster states [19, 8, 20, 21].
@@
The preparation of multiparticle complete graph states can be conceptually divided into two primary categories. The first approach is applicable to arbitrary graph states, employing the mathematically proven two-qubit controlled-Z gate (CZ) [22, 23, 24]. However, these protocols are constrained by the operational time and fidelity of the logic gate. On the contrary, complete multiparticle graph states can also be generated within the framework of cavity quantum electrodynamics [25, 26, 27]. In particular, certain theoretical schemes [28, 28, 29] enable one-step realization, but these one-step approaches require uniform interaction strengths between all particle pairs, which poses challenges for experimental implementation. Recently, the proposal of an equivalent Hamiltonian (Heisenberg $XXX$ model with a staggered field) [30] has sparked interest in Floquet engineering for the construction of such a Hamiltonian [31, 32]. Nevertheless, in practice, controlling the impact of noncommutation of the Hamiltonian under periodic driving becomes challenging as the particle number increases.
@@
As quantum control theory progresses, optimal quantum control provides a toolkit to craft external field shapes to perform assigned tasks, such as state generation [33, 34, 35, 36], quantum gate implementation [37, 38], and noisy intermediate-scale quantum technology [39]. These tools aim to achieve the specified objectives with minimal energy and resource expenditure. The practical realization of calculated field shapes has been facilitated by the availability of commercially accessible arbitrary waveform generators. In this paper, using the Gradient Ascent Pulse Engineering (GRAPE) optimization algorithm [40, 41], we propose an alternative Hamiltonian represented as a spin-$1/2$ Heisenberg $XX$ chain with identical nearest-neighbor interactions under an optimized time-dependent magnetic field. Governed by this Hamiltonian, we successfully generate complete graph states for a particle number ranging from $N=3$ to $N=6$. Our protocol boasts two key advantages: (i) It exclusively relies on the nearest-neighbor interactions, ensuring robust operability and scalability in experimental setups. (ii) The operational time is significantly reduced to less than $0.25~{}\mu{\rm s}$, a crucial factor in mitigating the effects of decoherence.
@@
The structure of this paper is organized as follows. In Sec. II, we initiate our exploration with the introduction of the definition of the complete graph state and an analysis of the corresponding scheme. Subsequently, in Sec. III, we employ the GRAPE optimization algorithm to articulate a system-independent realization of complete graph states in a “one-step” fashion. Transitioning to Sec. IV, we delve into a specific physical system, the Rydberg atom array, providing a comprehensive examination of the experiment’s feasibility. Our results suggest that, taking into account experimental noise, the population of the complete graph state with $N=6$ can exceed $0.91$ in a time frame of $0.25~{}\mu{\rm s}$. Finally, Sec. V encapsulates our findings, offering a concise summary.
@@
Figure 1: (a) Multiparticle complete graph states with the number of vertices $N=3\sim 6$. The graph connecting the two associated vertices has an edge whenever two particles have interacted. (b) Experimental geometry. Particles are arranged as a chain with the open boundary condition and have equivalent interactions between the nearest particles.
@@
with the convention $\sigma_{z}^{N+1}\equiv 1$. To generate such complete graph states, we construct a Hamiltonian that holds a global time-dependent magnetic field, which takes the form of ($\hbar=1$)
@@
$H_{\textrm{XX}}=\sum_{\langle i,j\rangle}\frac{J}{2}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j})$ (3)
@@
Since the two parts ($H_{\textrm{XX}}$ and the global magnetic field term) commute with each other, the evolution can take place under the Hamiltonian (2) or under the $H_{\textrm{XX}}$ and the global magnetic field separately. Theoretically, any equivalent $H_{\textrm{XX}}$ model can be applied to this scenario. Under a global magnetic field with constant amplitude, an exact analytical solution of $B$ can be found for $N=3$, as shown in Appendix A. However, for $N>3$, such an exact solution is non-existent. Thus, we consider a time-dependent magnetic field $B(t)$ and introduce the GRAPE optimization algorithm to find the optimal solution for the system. In this way, we can expect to obtain a multiparticle universal protocol.
@@ III The utilization of the GRAPE optimization algorithm
Given the known form of the target state $|K_{N}\rangle$, the optimization problem, which is implemented using the gradient descent algorithm, is translated into the search for the optimal amplitudes $B(t)$ of the magnetic field. These amplitudes aim to guide the initial state $\rho(0)$ to $\rho(T)$ with maximum similarity to $\rho_{K}=|K_{N}\rangle\langle K_{N}|$ within a specified time duration $T$. In our protocol, the system Hamiltonian is segmented into two distinct parts.
@@
where $H_{z}=\sum_{i=1}^{N}S^{z}_{i}$ is referred to as the control component. The quantum control landscape, depicting the expectation value of a Hermitian observable operator $\rho_{K}$ at a given time $T$, is represented as [42]
@@
Target state Evolution time (units of $J$) Population
@@
The results demonstrate the successful generation of multiparticle complete graph states under the influence of a spin-$1/2$ Heisenberg $XX$ chain in a time-varying magnetic field optimized by the GRAPE algorithm. Consequently, the preparation of multiparticle complete graph states can be achieved mathematically in a “one-step” manner through Eq. (2). This protocol is universally applicable in systems capable of nearest-neighbor interactions characteristic of the spin-$1/2$ Heisenberg $XX$ chain, including superconducting quantum circuits [43, 44, 45, 46] and ion traps [47, 48, 49, 50, 51]. In the following part, we will delve into the detailed methodology for preparing the aforementioned multiparticle complete graph states within a neutral-atom system, closely considering experimental parameters.
@@ IV A specific implementation based on Rydberg atom array
Figure 3: The realization of the multiparticle complete graph states with $N=3\sim 6$ governed by the Hamiltonian $H_{\textrm{sys}}$ initialized from state $|\Psi_{0}\rangle=\bigotimes_{i=1}^{N}[{1}/{\sqrt{2}}(|\uparrow\rangle+|\downarrow\rangle)]$. (a) and (b), respectively, show evolution under the magnetic field in Gaussian and random types for $N=3$. The dotted lines correspond to the evolution under the guess field, whereas the solid lines correspond to the optimized one. Then (c) and (d) show the evolution for $N=4$, while (e) and (f) for $N=5$, (g) and (h) for $N=6$. The insets show the corresponding time-dependent $B(t)$ of the initial guess and the optimized result, where $B_{0}=2\pi\times 1~{}{\rm MHz}$.
The Rydberg atom array, characterized by robust programmability, has recently emerged as a powerful candidate in quantum simulations, spurred by breakthroughs in experimental technology [52, 53]. Individual Rydberg atoms, for example, can be placed in an optical tweezers array [54, 55]. Leveraging the long-range interactions inherent in highly excited Rydberg states, the Rydberg atom array finds widespread applications in quantum state transmission [56], quantum topology research [57, 58, 59], and artificial gauge field [60, 61]. Notably, recent works, such as those in Refs. [62, 63], have realized a primary quantum computer employing 256 Rydberg atoms, demonstrating proficiency in various quantum tasks.
Within the Rydberg atom array, the Rydberg dipole-dipole interaction serves as a direct means to achieve $H_{\textrm{XX}}$ [64, 65]. Additionally, the magnetic field $B(t)$ can be equivalently obtained through the Stark effect induced by a large detuned time-dependent laser field coupled transition between the intermediate state $|p\rangle$ and one of the Rydberg states [66, 67]. It is crucial to note that the inevitable long-range interactions among Rydberg atoms, such as van der Waals (vdW) forces and the long-range dipolar interactions, should also be taken into consideration, with respect to the realistic physical system.
@@ IV.1 The realization of multiparticle complete graph states
We consider a one-dimensional chain geometry with open periodic boundary conditions, as shown in Fig. 1(b), in which $N$ atoms are contained in tweezers that are equidistantly spaced $R$ along the quantized $z$ axis. For the purpose of illustration, states with opposite parity are transformed into pseudo-spin states, using the ${}^{87}\textrm{Rb}$ atom as an example, $|\uparrow\rangle=|80S_{1/2},m_{j}=1/2\rangle$ and $|\downarrow\rangle=|79P_{3/2},m_{j}=1/2\rangle$ [68]. The resonant dipole-dipole interaction gives rise to the Heisenberg $XX$ Hamiltonian
$H_{\textrm{XX}}=\sum_{\langle i,j\rangle}\frac{V_{\textrm{dip}}(\theta)}{2}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}),$ (12)
where $V_{\textrm{dip}}(\theta)=C_{3}(1-3\cos^{2}\theta)/R^{3}$ with $\theta$ the polar angle between the quantization $z$ axis and the vector direction of atomic connection, and the constant $C_{3}$ is defined as $2\pi\times 8.780~{}{\rm GHz}\cdot\mu{\rm m}^{3}$ [69]. For $\theta=0$ and $R=19.3~{}\mu{\rm m}$, we have $V_{\textrm{dip}}\simeq-2\pi\times 2.443~{}{\rm MHz}$. Additionally, we account for inevitable interactions, including long-range vdW interactions and dipolar interactions, which are characterized as error terms.
$H_{\textrm{err}}$ $=$ $\sum_{j>i}(U^{\uparrow}_{ij}|\uparrow_{i}\uparrow_{j}\rangle\langle\uparrow_{i}\uparrow_{j}|+U^{\downarrow}_{ij}|\downarrow_{i}\downarrow_{j}\rangle\langle\downarrow_{i}\downarrow_{j}|)$ (13)
$+\sum_{j>i+1}\frac{V_{\textrm{dip}}^{ij}}{2}(\sigma^{x}_{i}\sigma^{x}_{j}+\sigma^{y}_{i}\sigma^{y}_{j}),$
where $U^{\uparrow(\downarrow)}_{ij}=-C^{\uparrow(\downarrow)}_{6}/R_{ij}^{6}$ with $C^{\uparrow}_{6}=-2\pi\times 4161.55~{}{\rm GHz}\cdot\mu{\rm m}^{6}$ and $C^{\downarrow}_{6}=2\pi\times 3452.60~{}{\rm GHz}\cdot\mu{\rm m}^{6}$, and $V_{\rm dip}^{ij}=C_{3}(1-3\cos^{2}\theta)/R_{ij}^{3}$. $R_{ij}$ denotes the distance between the $i$th and $j$th atom.
@@
Using $N=3\sim 6$ as illustrative examples, we depict the evolutions governed by Eq. (14). In Fig. 3, we specifically show the evolutions under the initial guess and optimized fields in Gaussian and random types. Within this framework, the populations in the target state, denoted as $P(N)_{O}$, exhibit multiple peak points corresponding to different values of $T$ (as shown in the picture in Appendix B). By scanning the time within $0.25~{}\mu{\rm s}$, we successfully obtain the optimized fields under $H_{\textrm{sys}}$ for various numbers of particles. The inset shows the time-dependent $B(t)$ of the application, where solid lines represent the optimized pulse, and dotted lines represent the guess pulse. The evolution time $T$ and the populations of the target states are summarized in Table 2. Although the population of the target state decreases as $N$ increases, we can still prepare the six-particle complete state with a population greater than $0.92$.
@@
$|K_{3}\rangle$ 0.141 0.9989
$|K_{4}\rangle$ 0.172 0.9920
$|K_{5}\rangle$ 0.203 0.9728
$|K_{6}\rangle$ 0.233 0.9294
Recently, a quantum circuit utilizing the Rydberg atom array has been demonstrated to generate a complete graph state of five particles [70]. When coupled with quantum annealing algorithms, the fidelity of the target state can be increased from approximately $0.85$ to $0.999$ by increasing computational depth. However, this increased computational depth is accompanied by prolonged execution times, leading to an increased susceptibility to decoherence effects. On the contrary, our protocol enables the generation of the target state in a single step, regardless of the number of atoms involved. This approach significantly reduces operational time while preserving high levels of fidelity, which presents notable advantages in mitigating decoherence effects.